OpenStudy (anonymous):

Can you give me an example how to solve -6=-9/x

jhonyy9 (jhonyy9):

3/x =2/5 cross multiplie 3*5=2*x 15=2x divide both sides by 2 15/2 =x

OpenStudy (bittuaryan):

\[-6=\frac{ -9 }{x }\]\[\rightarrow -6x=-9\]or \[6x=9\]\[x=\frac{ 9 }{ 6 }\]\[\rightarrow x=\frac{ 3}{ 2 } \] or 1.5